40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 19 21:10:27 2010, yuta, Update, CDS, bad filter coefficients for every suspensions! 
    Reply  Wed Oct 20 01:22:18 2010, yuta, Update, CDS, fixed filters for MCs, but no damping Screenshot.png
Message ID: 3744     Entry time: Wed Oct 20 01:22:18 2010     In reply to: 3742
Author: yuta 
Type: Update 
Category: CDS 
Subject: fixed filters for MCs, but no damping 

(Rana, Yuta)

Summary:

 The damping servo for MCs has not been working since Monday.
 We already found that some filter coefficients were wrong.(see elog #3742)
 So, we fixed it (just for MCs now).
 But still doesn't work.

What we did:
Fixed the filters for MC suspensions;
 1. In /cvs/cds/rtcds/caltech/c1/chans/ directory, we replaced C1MCS.txt with ./filter_archive/c1mcs/C1MCS_101019_101927.txt.
  This is the archive of the filter bank for MC suspensions, when filter designs were correct(when we hadn't opened it with foton yet).

 2. Deleted all the filter coefficients.
   By using emacs magical C-<space> C-x r k.

 3. Loaded with foton to get correct coefficients.

 4. Reloaded coefficients for C1MCS.

We now have the correct filters, but the damping still doesn't work.

 5. To confirm that the filters are now correct and the model is working correctly, we measured the transfer function between ULSEN input and ULCOIL output and compared with the calculated TF from the filter bank file (Attachment #1) .

 6. To calculate the designed function, we used the following matlab scripts;
    /cvs/cds/caltech/users/rana/mat/utilities/FotonFilter.m        # takes the foton file and returns the TF
    /cvs/cds/caltech/users/rana/mat/utilities/readFilterFile.m    # reads the foton file
    /cvs/cds/caltech/users/yuta/scripts/sentocoil.m                   # calculates the total TF from SEN to COIL

Result:
 As you can see from the attachment, filters seem fine now. (The red line is the measured and the blue line is the calculated function in the plot)
 But the damping (for POS, PIT, YAW) still doesn't work. Why????

Next work:

 - see if coils are pushing correctly
 - push cable connectors further!
 - check input and output filters (whitening, dewhitening)
 - when fixed, use my fully updated QAdjuster.py for adjusting Q for multiple channels automatically!!

Attachment 1: Screenshot.png  66 kB  Uploaded Wed Oct 20 02:25:30 2010  | Hide | Hide all
Screenshot.png
ELOG V3.1.3-